OpenStudy (anonymous):

Evaluate the integral. 2√(x+1)dx

OpenStudy (anonymous):

hint \[\sqrt{x+1}=(x+1)^{\frac{1}{2}}\] use the power rule backwards

OpenStudy (mathmale):

Hint #2: It may help to use the substitution u = x + 1. Then du = dx. What next?